# Generative Logic: A New Computer Architecture for Deterministic Reasoning and Knowledge Generation

### v3

@@ 6.5 The FTA Campaign
Once the above stack is complete, the FTA campaign becomes a direct engineering effort. A route of auxiliary facts and intermediate theorems leading from the Peano axioms toward FTA will be identified and pressure-tested in sequence. If the conjecturer fails to surface a required fact, the conjecturer is extended. If the prover cannot establish it, the prover is extended. That loop repeats until the route is carried end to end.
@@ 6.6 Path to Scalability
In parallel with the mathematical extensions above, the implementation will be progressively hardened. The Python conjecturer — currently the dominant runtime cost — will be migrated to C++, recovering the multithreading headroom that Python’s GIL forecloses. The architecture will then be adapted for massively parallel cloud execution, distributing the workload across large numbers of instances on platforms such as AWS or GCP. The long-term target substrate remains a cloud of specialized ASIC chips, where the full parallelism of the Logic Block grid can be realized without constraint.
6.7 LLM Deployment
LLM Interface & Feedback Loop (post-ENT). After the ENT milestone, we will expose tool calls so an LLM can (i) submit MPL definitions and conjectures, (ii) invoke GL runs, and (iii) inspect derived theorems and proof graphs. In the reverse direction, GL becomes an LLM reasoning booster: when the model confronts a mathematical (and eventually any MPL-encodable) reasoning task, it can translate that task to MPL and ask GL to deterministically study its deductive consequences, feeding validated results back into the dialogue. GL remains an external proof engine — not a Python library import — so the LLM/GL handshake mirrors, but does not collapse into, standard Python execution flows.
@@ 8 Conclusion
The implementation is a prototype. Its performance figures are those of a prototype. The path to larger theories is identified and involves no steps that appear, at present, to be impossible.
@@ 9 List of Abbreviations
ENT
Elementary Number Theory
FTA
Fundamental Theorem of Arithmetic
AWS
Amazon Web Service
GCP
Google Cloud Platform
CE
Counterexample
NoC
Network-on-Chip
SMT
Satisfiability Modulo Theories

### v4

@@ 7.2 The FTA Campaign
The FTA campaign is a direct engineering effort. A route of auxiliary facts and intermediate theorems leading from the Peano axioms toward FTA will be identified and pressure-tested in sequence. If the conjecturer fails to surface a required fact, the conjecturer is extended. If the prover cannot establish it, the prover is extended. That loop repeats until the route is carried end to end.
7.3 Toward a Full-Provenance CAS
After FTA, the next major target is algebraic number theory, pursued in an agentic loop with a code-capable LLM such as Claude Code (CC). The first step is to establish a $p$-adic number formalism within MPL, beginning the transition from ENT to Algebraic Number Theory (ANT). The natural numbers 0 through $n$ will be derived and handled exactly as they are now in the Incubator; from this base, the $p$-adic number system will be introduced with all its necessary properties auto-derived by GL in the usual manner.
The full CAS will begin with integers: once integer arithmetic is established with full provenance, the system will extend to fixed-point and floating-point representations. The long-term goal — and decades-old vision of the GL project — is to produce a calculator in which not just the method of calculation is founded in axioms, but every single digit of the output carries a full provenance graph, traceable from the final result back to the foundational definitions.
@@ 7.4 Path to Scalability
The architecture will be adapted for massively parallel cloud execution, distributing the workload across large numbers of instances on platforms such as AWS or GCP. The long-term target substrate remains a cloud of specialized ASIC chips, where the full parallelism of the Logic Block grid can be realized without constraint.
7.5 LLM Integration
A first trial of CC driving GL — reading existing MPL definitions, formulating new ones in valid MPL syntax, and writing configuration files — is documented in a publicly available video recording. [^1] https://www.youtube.com/watch?v=oJmBJNc1xVA While early-stage and unspectacular in its scope, this trial answers the most important question unequivocally: an LLM can use GL.
The main bottleneck at this stage is not LLM capability but GL maturity and the absence of a persistent harness that allows CC to remain in an iterative loop until a task is solved — inspecting intermediate results, adjusting definitions or configurations, and re-running the pipeline as needed. Building this harness is an engineering task; the scientific question of whether an LLM can interface with a deterministic reasoning engine is settled by the demonstration itself.
In the reverse direction, GL becomes an LLM reasoning booster: when the model confronts a mathematical (and eventually any MPL-encodable) reasoning task, it can translate that task to MPL and ask GL to deterministically study its deductive consequences, feeding validated results back into the dialogue. GL remains an external proof engine — not a Python library import — so the LLM/GL handshake mirrors, but does not collapse into, standard Python execution flows.
@@ 9 Conclusion
Beyond proof generation, the architecture includes three additional pipeline stages. The Incubator auto-generates the ground-level fact tables that feed the CE filter, closing the bootstrap loop from definitions to ground facts without requiring case differentiation. The Compressor eliminates post-proof redundancy, reducing each batch’s theorem mass to a minimal essential core. The Verifier independently checks every inference step in the processed proof graph, confirming zero failures across 34,320 checks. Together, these stages make the pipeline self-sufficient and externally verifiable.
@@
The implementation is a prototype. The path to larger theories is identified and involves no steps that appear, at present, to be impossible.
@@ 10 List of Abbreviations
AWS
Amazon Web Service
CAS
Computer Algebra System
$C$
Claude Code
CE
Counterexample
ENT
Elementary Number Theory
FTA
Fundamental Theorem of Arithmetic
GCP
Google Cloud Platform
@@
NoC
Network-on-Chip
RT
Run Time
SMT
Satisfiability Modulo Theories